Solve the initial value problem.

2y” - 11y’ + 12y =0 ; where y(0) = 5 and y’(0) = 15


1
Expert's answer
2021-09-05T16:02:22-0400

Given, "2y''-11y'+12y=0."

Auxiliary equation,

"2m^2-11m+12=0\\
(m-4)(m-\frac{3}{2})=0\\
m=4,\frac{3}{2}"

Therefore, general solution is

"y(x)=c_1e^{4x}+c_2e^{\frac{3}{2}x}.\\
y'=4c_1e^{4x}+\frac{3}{2}c_2e^{\frac{3}{2}x}"

Given, the initial condition "y(0)=5 \space and \space y'(0)=-15" .

"y(0)=c_1+c_2\\
\implies 5=c_1+c_2\\
and\\
y'(0)=4c_1+\frac{3}{2}c_2 \\
\implies -15=
4c_1+\frac{3}{2}c_2"

Solving above two equations, we get

"c_1=-9, c_2=14"

Therefore, particular solution of the given IVP is "y(x)=-9e^{4x}+14e^{\frac{3}{2}x}" .